In 2018, when Noah Corporation was organized, it immediately issued 790 shares of $8 par, 7 percent, noncumulative preferred stock and 1,563 shares of $2 par common stock. It is now 2020 and the company is declaring its 1st dividend. No dividends were declared in 2018 or 2019.

Assume that the board of directors declares a $5,800 cash dividend at the end of 2020. How much of the dividend will go to the common stockholders? $______
Answer #1

Dividend to preferred stock holders = 790*$8*7% = $442.40

Dividend to the common stock holders = 5800-442.40 = $5357.60